OpenStudy (anonymous):

\[\sqrt[8]{x-4}*\sqrt[8]{(x-4)^{6}}\]

OpenStudy (anonymous):

(x-4)^ 1/8 * (x-4)^ 6/8 = (x-4)^ 1/8+6/8 = (x-4)^7/8